2 trains - 30 kmph each, headed towards each other on same track.
A bird ( v = 60 kmph ) flies off the front of 1 train when they r 60 m  apart & heads directly towards other train & so forth. Total distance the bird travels ???

Grade:12

1 Answers

Puneet Mittal AskiitiansExpert-IITD
13 years ago

Hi Mohit,

The Answer is ,

Let the distance between the trains be x km Initially.

The relative speed of trains is 30 + 30 = 60 kmph.

Thus distance x is reducing with speed 60 kmph.

Now t = total time taken when the trains strike each other

Thus t = x / 60 hrs. where x is in kms.

Now the distance travelled by bird is S = speed of bird and time the bird is alive.

Thus , S = 60 * t

            = 60 * x / 60

            = x  kms.

       If Initially the trains are 60 m apart then total distance the bird will travel is 0.06 kms

      of if Initially the trains are 60 km apart then total distance the bird wiil travle is 60 km.

Please feel free to ask as many questions u have.

Puneet